What is the instantaneous acceleration at t = 2.1 s of a particle whose position is given by y(t) = -4.1t2 + 6.7?




1
Expert's answer
2021-10-05T15:42:40-0400

Let's first find the velocity of the particle:


"v(t)=\dfrac{dy(t)}{dt}=-8.2t."

Then, we can find the acceleration of the particle:


"a(t)=\dfrac{dv(t)}{dt}=-8.2\ \dfrac{m}{s^2}."

Since the expression for the acceleration don't depend on "t" , the acceleration of particle for each "t"

will be "a(t)=-8.2\ \dfrac{m}{s^2}." Therefore, the instantaneous acceleration at t = 2.1 s equals "-8.2\ \dfrac{m}{s^2}".